### v1

# V-LASIK: Consistent Glasses-Removal from Videos Using Synthetic Data

Abstract.
Diffusion-based generative models have recently shown remarkable image and video editing capabilities. However, local video editing, particularly removal of small attributes like glasses, remains a challenge. Existing methods either alter the videos excessively, generate unrealistic artifacts, or fail to perform the requested edit consistently throughout the video.
In this work, we focus on consistent and identity-preserving removal of glasses in videos, using it as a case study for consistent local attribute removal in videos. Due to the lack of paired data, we adopt a weakly supervised approach and generate synthetic imperfect data, using an adjusted pretrained diffusion model. We show that despite data imperfection, by learning from our generated data and leveraging the prior of pretrained diffusion models, our model is able to perform the desired edit consistently while preserving the original video content. Furthermore, we exemplify the generalization ability of our method to other local video editing tasks by applying it successfully to facial sticker-removal. Our approach demonstrates significant improvement over existing methods, showcasing the potential of leveraging synthetic data and strong video priors for local video editing tasks.
Figure 1. Our method receives an input video of a person wearing glasses, and consistently removes the glasses, while preserving the original content and identity of that person. Our method successfully removes the glasses even when there are reflections (bottom-middle example), heavy makeup (top-right), and eye blinks (bottom-right). Red rectangles are zoomed-in at the bottom row.
[…]
1. Introduction
Recent advances in diffusion-based generative models (Ho et al., 2020; Sohl-Dickstein et al., 2015) have demonstrated impressive capabilities in image and video editing (Rombach et al., 2022; Zhang et al., 2023a; Geyer et al., 2023; Yang et al., 2023; Zhang et al., 2023b; Kara et al., 2023; Bar-Tal et al., 2024). While local editing, such as removing and adding attributes without changing the rest of the content, mainly works for images, local video editing remains a challenge. Video frames often contain motion blur and challenging poses that are less common in images. Moreover, videos of people pose additional challenges, as humans are highly sensitive to subtle unrealistic “uncanny” artifacts, and our visual system is more sensitive to motion inconsistencies, so the quality bar is higher. Existing image editing and inpainting methods (Nitzan et al., 2024; Zhang et al., 2023a; Tsaban and Passos, 2023; Brooks et al., 2023; Couairon et al., 2023; Meng et al., 2021) are trained on clean images, that do not contain the aforementioned problems, and therefore do not perform well over video frames; They typically either change the original frames too much, or do not perform the requested change completely, as illustrated in Fig. 2.
Video inpainting methods (Zhou et al., 2023; Yu et al., 2023a; Li et al., 2022b; Zhang et al., 2022) typically remove objects by deleting and filling regions of the video, often disregarding the information originally present in those regions. When it comes to generating new video content behind an object that never moves, these methods tend to generate a low detailed background, or one with clear unrealistic artifacts. As a representing use-case, we choose to tackle glasses removal, as it is a particularly challenging example; where some content is never or only partially revealed (e.g. eyes, eyebrows), and is changed across the video (e.g. eyelids change position across frames). Since eyes contribute greatly to identity perception, a consistent and highly realistic generation is required for an identity to be preserved. However, current video inpainting works typically generate smooth results when using small masks, that do not fit the facial case well (see ProPainter (Zhou et al., 2023) and FGT (Zhang et al., 2022) results in Fig. 5).
Mask-free video editing methods (Geyer et al., 2023; Zhang et al., 2023b; Khachatryan et al., 2023; Kara et al., 2023; Cong et al., 2024) mainly focus on changing the style of the whole video, or changing the appearance of large parts of it, and not on local changes that should adhere to the original context. Other works (Kasten et al., 2021; Bar-Tal et al., 2022) learn an atlas for each new video, leading to long inference time. Furthermore, as mentioned in Text2Live (Bar-Tal et al., 2022), the goal of these works is to manipulate the appearance of existing objects, not to change the composition of the elements in the scene. Hence, their method is not capable of successfully removing attributes such as glasses from faces.
In this work, we focus on removing glasses from videos as a case study for local video editing. We introduce a new method that performs this task consistently while preserving the original identity. To the best of our knowledge, no paired data exists for this task; Therefore, we cannot finetune a model directly to solve it. To tackle this problem, we generate imperfect synthetic data, and use it to train a model in a weakly supervised manner. We show that by using an adjusted pretrained image inpainting model, we can generate imperfect data pairs, and learn from them. Our method is closest to video inpainting, in the sense that it removes something from the video and replaces it with different content. However, similarly to recent video editing works, the input to our model is a mask-free image, allowing the model to use facial details that are occluded for inpainting models. As glasses often contain colored lens and reflections that change across frames, which affect the way the eye behind them looks, they are a challenging attribute to remove from faces. Hence, we incorporate cross-frame attention layers in our data generation model, which allows aggregating information from different frames. Still, as presented in Fig. 2, the generated data (‘Synth data’) is imperfect and contains many artifacts, such as not always preserving the eyelid position and other features from the original frame. To overcome these challenges, we finetune a diffusion based image-to-image model over this data. This way we can leverage the prior of the pretrained model, while enhancing its capabilities with our new data. Additionally, the non-masked input allows the model to exploit the fact that it is exposed to useful information within the masked region to achieve the requested edit, while preserving the rest of the content from the original frames. Despite the imperfections in our synthetic data, a model with a strong prior can leverage its prior knowledge to outperform the training data and produce high-quality results. Finally, to get temporally consistent results, we combine our trained model with a motion prior, and achieve state-of-the-art results on video glasses-removal, surpassing all current video editing and inpainting methods. To demonstrate that our method can be generalized to other local video editing tasks, we successfully apply it to the task of removing stickers from faces. We release code and both the stickers and the glasses removal datasets for further research.
[…]
Figure 2. Glasses-removal from a blinking eye by image editing methods Left to right: LEDITS (2023), Instruct pix2pix (2023), Lyu et al. (2022), Stable Diffusion inpaint (2022), ControlNet inpaint (2023a), our synthetic dataset generation result, and our final result. As image editing methods expect high quality images with people looking straight to the camera, they struggle when these constraints are not met. In our dataset result ‘Synth data’, as a result of the cross-frame attention, eye artifacts appear. However, our model is still able to learn from the imperfect data and remove the glasses better than any out-of-the-box method, and better than the data it was trained on.
2. Related Work
2.1. Image editing
Since the advent of diffusion models for image generation and editing (Ho et al., 2020), real image editing has rapidly advanced and showed remarkable results (Brooks et al., 2023; Couairon et al., 2023; Tsaban and Passos, 2023; Hertz et al., 2022; Tumanyan et al., 2023; Zhang et al., 2023a; Meng et al., 2021), including local editing such as removing glasses from images. Moreover, image inpainting methods (Nitzan et al., 2024; Rombach et al., 2022; Zhang et al., 2023a; Yildirim et al., 2023), have also shown impressive results, realistically replacing the content behind a given mask. Additionally, prior work (Lyu et al., 2022; Lee and Lai, 2020) tried to tackle the task of removing glasses from images directly. Indeed, these works perform well over images. However, as shown in Fig. 2 they fail when applied to video frames of people, that do not always look directly at the camera, and constantly move, causing motion blur and other artifacts. Moreover, when applied frame-by-frame, the generated content differs between frames, resulting in temporal inconsistencies.
2.2. Video editing
Recently, video editing has also developed greatly with video-to-video translation methods (Geyer et al., 2023; Khachatryan et al., 2023; Cong et al., 2024; Kara et al., 2023; Qi et al., 2023; Wu et al., 2023; Zhang et al., 2023b) that focus on transforming the entire frame into a different style, while trying to preserve temporal consistency in the generated videos. Some of them, such as RAVE (Kara et al., 2023), or TokenFlow (Geyer et al., 2023) with SDEdit (Meng et al., 2021) also perform local editing. However, in the case of glasses-removal, where a person moves across the video and the requested edit is small and delicate, many artifacts such as face deformations and inconsistencies are generated. Examples are presented in Fig. 5 and in video results in the supplementary material. Another line of works is atlas-based video editing (Kasten et al., 2021; Lu et al., 2021; Bar-Tal et al., 2022; Suhail et al., 2023). These methods learn an atlas for each video, to apply changes to either the background or foreground of the video. The initial atlas reconstruction requires excessive computational resources and long running times. Moreover, as these methods were designed to change the appearance of existing objects, they are not meant for adding or removing attributes that were not part of the original video, such as glasses, and do not perform these changes well. Additionally, while theoretically allowing to remove layers from a video, these methods are limited to relatively simple motion and to static backgrounds.
2.3. Video inpainting
Video inpainting is the task of consistently filling-in new content behind a given mask throughout a video (Yu et al., 2023a; Zhang et al., 2022; Zhou et al., 2023; Gu et al., 2023; Li et al., 2022b). Such methods work well when the object moves throughout the video, and the model is able to fill-in the background from other video frames, where the background is visible. However, they struggle with filling-in completely occluded areas, especially when they are part of the foreground, such as faces. Particularly, areas behind glasses are only partially revealed by head motion, and sometimes glasses cause optical deformations that hide the background. Slight changes to face features in the “background” of glasses may yield unrealistic results, or alter the face identity, as shown in the ProPainter (Zhou et al., 2023) and FGT (Zhang et al., 2022) results in Fig. 5. While such changes to smooth backgrounds may not significantly affect the perceptual quality of the results, in foreground areas and particularly in faces they are detrimental.
2.4. Learning from synthesized data
Deep learning models require large amount of data to learn from, however paired data is not always available for the task at hand. Hence, several works (Li et al., 2022a; Peebles et al., 2022; Ravuri and Vinyals, 2019; Brooks et al., 2023; Lyu et al., 2022) use generative models for data generation to train models. These works usually rely on the high quality of the generated data, and learn solely from it. In this work, we accept that generated data is usually imperfect, and is insufficient to achieve the required result on its own, thus we take advantage of the strong priors of pretrained models and use our generated data for fine-tuning. This way, the model learns to generate results that are superior to the data it was trained on, while performing the relevant task.
Lyu et al. (Lyu et al., 2022) specifically explores synthetic data for the task of glasses-removal from images. However, their data acquisition process requires face scanning and 3D data, which is hard to acquire, unlike our method which does not require any special equipment or effort.
3. Method
Input Masked input w/o CF attn w/ CF attn Input Masked input w/o CF attn w/ CF attn Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption input inaccurate mask glasses remnants no remnants input masked input reflections no reflections
Figure 3. Cross-frame attention importance in data generation. Cross-frame attention helps removing glasses remnants, even when the mask is not perfect (left example) and reducing glasses reflections (right example).
[…]
Figure 4. Method overview: Step 1: we create an imperfect synthetic paired dataset by generating glasses masks for each video frame and inpainting it. We inpaint each frame using an adjusted version ControlNet inpaint (Zhang et al., 2023a). We replace the self-attention layers with cross-frame attention (cf attn) and use blending between the generated latent images and the noised masked original latent images at each diffusion step. The generated data in the first step is imperfect; e.g. in the middle frame, the person blinks, however its generated pair has open eyes. Nevertheless, the data is good enough for finetuning an image-to-image diffusion model and achieving satisfactory results, due to the strong prior of the model. Step 2: Given our trained model for the task of removing glasses from images, we incorporate it with a motion prior module to generate temporally consistent videos without glasses from previously unseen videos. To obtain the original frame colors, at each diffusion step we blend the generated frames with the noised original masked latent images, and before decoding, we apply an Inside-Out Normalization (ION), to better align the statistics within the masked area and the area outside of the mask.
[…]
Given an input video of a person wearing glasses, our goal is to remove the glasses while preserving all other information. As illustrated in Fig. 4, our approach to removing glasses from videos consists of three stages: data generation, training, and editing. First, as no paired data is available for this task, we generate a synthetic paired dataset using videos of people wearing glasses. Next, we finetune an image-to-image diffusion model over our synthetic dataset, to get realistic video frames without glasses, where all other parts of the frame remain similar to the original frame, and the identity of the edited person is preserved. Finally, we incorporate our trained model with a pretrained motion module, into a video editing pipeline, to obtain temporally consistent videos without glasses.
3.1. Paired Data Generation
As illustrated in Fig. 2, current image editing and inpainting methods do not perform well on the task of removing glasses from video frames. Therefore, we wish to train a model for this task. However, to the best of our knowledge, no relevant paired dataset currently exists. Hence, we create a synthetic dataset of paired video frames, with and without glasses, using videos from the CelebV-Text dataset (Yu et al., 2023b) where people wear glasses. For each video frame, we generate a glasses mask using a face parser (Zheng et al., 2022) that identifies the glasses in it. The eyes are a key component in the identity of a person, and the current position of the eyelids is crucial to generating a result that is consistent with the original video, that might contain eye-closure and blinking. Therefore, we want to give the model information about the eyes and eyelids, to keep their original appearance. To do that, we make eye “holes” in the glasses masks, using the identified eye landmarks. Then, we inpaint the glasses area by applying an adjusted inpainting diffusion model over each video frame and its generated mask.
We make two adjustments to the inpainting model: First, to achieve smoother and more realistic transitions between the frame and the inpainted part, we blend the latent feature vectors with a noised encoded version of the masked original frame at each diffusion step, as suggested in Blended Latent Diffusion (Avrahami et al., 2023). Second, to increase the edit consistency across frames from the same video, we replace the self-attention layers of the model with a cross-frame attention (Khachatryan et al., 2023). For each video frame, we perform a cross-frame attention with $k$ reference frames from the same video. We use multiple reference frames as sometimes information such as the eyebrows or the eye color is hidden behind the glasses or their reflections. Using cross-frame attention with multiple frames allows for generalization from all $k$ frames. For this purpose, we adopt the cross-frame attention mechanism suggested by TokenFlow (Geyer et al., 2023):
(1) $\text{softmax}(\frac{Q\cdot[K_{0},...,K_{k}]^{T}}{\sqrt{d}})\cdot[V_{1},...,V_{k}]$
where each query frame ($Q$) attends to $k$ different values ($V_{i}$), corresponding to $k$ different reference frames in our case. As a result, even when face parts are occluded, the information about the occluded parts can be retrieved from the reference frames if they are revealed there. An example of occluding reflections is presented on the right side of Fig. 3. In this example, there are reflections over the glasses, hiding the true eye colors behind them. Without using cross-frame attention, reflections and artifacts are present on the output result. However, when using cross-frame attention, the true eye color is preserved and no reflections are shown in the result. Moreover, using cross-frame attention also helps when the masks are not exact and do not cover all of the glasses in all video frames, as in the left example of Fig. 3. In this case, glasses remnants appear when no cross-frame attention is used, due to an incomplete mask. However, they are removed when we use cross-frame attention, where the masks of the reference frames cover the uncovered part.
3.2. Model fine tuning
After the data generation stage, we obtain pairs of frames with and without glasses. However, our generated frames are not temporally consistent, and also contain per-frame artifacts in many cases, as featured in Figs. 2 and 4. For example, the eyelid positions, i.e. closed or open, sometimes change after the inpainting process. Despite that, the generated data is good enough for finetuning a pretrained image-to-image diffusion model for the task of removing glasses from faces. These models have strong prior for generating realistic looking images that are similar to the original ones and preserve their small and delicate details. Therefore, after finetuning, the model learns the task of removing glasses from our data, while preserving fine details such as eye color and eyelid positions.
3.3. Video editing pipeline
To consistently remove glasses from unseen videos, we integrate our trained model with a pretrained motion module (Guo et al., 2024). When applied directly to video frames, this model’s results often exhibit slightly different colors from the original input. To overcome this issue, we can use the masks from the data generation stage to retrieve the original values on areas that are outside of the mask. For a smooth result, at each diffusion step we perform a gradual blending between the noised masked input and the generated result, such that the area within the mask changes completely, and the areas that surround the mask change less, gradually decreasing the amount of change as the pixels are further from the mask. Additionally, inspired by AdaIN (Huang and Belongie, 2017), we apply a new normalization function, we dub Inside-Out Normalization (ION), where we aim to align the statistics of the masked area with those of the non-masked area. Formally, we calculate the mean and standard deviation of the masked area and the non-masked area, $\mu_{m},\sigma_{m}$ and $\mu_{\bar{m}},\sigma_{\bar{m}}$ respectively, and we normalize the values of the latent features in the area inside the mask by calculating:
(2) $ION(x)=\sigma_{\bar{m}}\frac{x-\mu_{m}}{\sigma_{m}}+\mu_{\bar{m}}$
ION allows a smooth transition between the areas inside and outside of the mask by moving the statistics of the latent masked area toward those of the non-masked area.
4. Implementation Details
[…]
Editing pipeline specifics. We use the pretrained motion module of AnimateDiff (Guo et al., 2024) with context length 16, context overlap 4 as our motion prior module. As shown in the right example of Fig. 9, the motion module tends to smooth the frames to get a more temporally consistent result, hence sometimes the results using this module get blurry. To avoid blurriness, we use only some of the motion layers and not all of them. Specifically, we remove the first 5 output motion layers. This way, we get a more realistic and less blurry video. For blending we use gradual values between 0 and 0.7 as mask values, so that pixels outside of the mask can also change a little bit for a smoother result.
Original colors vs. glasses removal trade-off. As an option in our video editing pipeline, we blend original pixels back into areas outside the masked region. If we do not blend them back, the colors in the edited video may not match those of the original video exactly. However, as the masks are not perfect, glasses remnants are sometimes left within the non-masked region, causing glasses remnants to appear in the result when using masks. To avoid it, we use dilation and different blending values outside of the mask. If the original colors are not as important and glasses-removal is of higher priority, one can use higher dilation for the mask, lower mask blending values, or even not use the masks at all.
General. We use Stable Diffusion 1.5 (rombach2021highresolution) as our backbone in all the method steps, as at the time of development there was no compatible motion module for SDXL.
5. Experiments
Input Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
T2V-Zero Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
TokenFlow Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
RAVE Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
CN inpaint Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
FGT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
ProPainter Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5. Visual comparisons: We compare our results to different video editing and inpainting methods. Other methods often struggle with glasses-removal, and even when they do remove the glasses, they tend to leave glasses remnants (e.g. RAVE right example), generate artifacts (e.g. FGT, ProPainter examples, TokenFlow left example), do not preserve the identity of the person (e.g. RAVE Left example), or their eyelids position (e.g. RAVE both examples).
[…]
Input Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
T2V-Zero Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
TokenFlow Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
RAVE Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
CN inpaint Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
FGT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
ProPainter Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6. Visual comparison: We compare our results to different video editing and inpainting methods. Other methods often struggle with glasses-removal, and even when they do remove the glasses, they tend to either change the identity completely (e.g. TokenFlow right example), generate artifacts such as black areas around the eyes (e.g. FGT, ProPainter both examples), or do not preserve the eyelids position (e.g. TokenFlow, RAVE right example)
[…]
We evaluate our results both qualitatively and quantitatively, testing three aspects:
1. Fidelity to the required edit, i.e glasses-removal.
2. Identity and content preservation of the original video. We want to remove the glasses while leaving the rest of the video intact.
3. Realism of the result, by means of temporal consistency and realism per-frame.
As we perform local video editing, we compare our results with the results of the SOTA video editing methods: TokenFlow (Geyer et al., 2023), RAVE (Kara et al., 2023), and Text2Video-Zero (Khachatryan et al., 2023). TokenFlow and Text2Video-Zero incorporate SDEdit (Meng et al., 2021) and instructPix2Pix (Brooks et al., 2023), respectively, to allow for local attribute editing, and RAVE uses CN “depth-zoe” (Zhang et al., 2023a) for that purpose. We also compare our model to the SOTA video inpainting works ProPainter (Zhou et al., 2023) and FGT (Zhang et al., 2022) as our work is similar to inpainting works in the sense that it tries to replace some part of the video. Additionally, we compare our results to the results of our video editing pipeline with CN inpaint (Zhang et al., 2023a) instead of our trained model, to emphasize that existing image inpainting models do not perform well enough on this task, even when combined in our video editing pipeline. Our quantitative and qualitative evaluations, including a comprehensive user study, demonstrate that our results are favored over all other methods across all tested aspects.
5.1. Experimental details
We generate our data pairs as described in Sec. 3.1 from the dataset CelebV-Text (Yu et al., 2023b). We train our model over 1296 of those videos, and test over 144 unseen videos.
5.2. Qualitative evaluation
Figure 7. User study results: First, we ask users which models remove the glasses from the input video (top). Then, when both models remove the glasses, we ask which one better preserves the identity of the person, which one is more realistic, and which result has less remnants of glasses (bottom). Compared to all model examined, the users preferred our results through all measured aspects.
[…]
Visual results of our method are presented in Figs. 5 and 6 and in the supplementary material. As illustrated in the figures, existing video editing and inpainting methods struggle with performing the required local edit, i.e., removing the glasses of the person. Moreover, even when they do remove the glasses, they tend to generate artifacts and unrealistic results, leave glasses remnants, or do not preserve the identity of the person or its original eyelid positions. In contrast, our method successfully removes the glasses while preserving the identity and content of the original video.
We perform a thorough user study to compare our results with the results of TokenFlow, RAVE, ProPainter, and our editing pipeline with CN inpaint. In the user study, we do not compare our results to those of Text2Video-Zero, as it only removes the glasses from the input video in 4% of the test videos, as featured in Fig. 5.
The user study tests three aspects: glasses-removal, resemblance to the identity in the original video, and realism of the result. For each video in the survey, we ask the users in which videos are the glasses removed. If both models removed the glasses for that video, we additionally ask which video contains less remnants of glasses, which one looks more realistic, and which one better preserves the identity of the person in the original video. We ask for remnants of glasses, because we noted that even when other models remove the glasses from the video, they often leave parts of the glasses or their reflections, e.g. as presented in the results of RAVE in Fig. 5. The user study contained side-by-side video pairs of our results vs. the results of different models for the same input (11 per model, 44 pairs in total), and it was answered by 57 users. The results of our user study are presented in Fig. 7. The top table shows the percentage of videos for which the users thought each model removed the glasses from the input video. The users thought our model removed the glasses in 98.9% of the cases, more than any other model. The bottom table shows the percentage of users that preferred our results over the tested models. Fig. 7 shows that the users preferred our results over all other models, in all the measured aspects: identity preservation, realism, and quality of glasses-removal.
5.3. Quantitative evaluation
Table 1. Quantitative results: We compare our results to different video inpainting and editing methods, where CN inpaint is CN inpaint embedded in our video editing pipeline instead of our model. We present results for two versions of our model — with and without masks, as elaborated in Sec. 5.3. We test the fidelity of the results by checking $\Delta G$, the average difference between glasses pixels in the original video frames and the edited one. We test the identity preservation of the edited video using $ID$ score, and the tradeoff between them using $ID\cdot\Delta G$. Moreover, we test the temporal consistency of the generated videos using the optical flow warp error $E_{warp}$.
Method $\Delta G_{(\times.01)}\uparrow$ $ID_{(\times.1)}\uparrow$ $ID\cdot\Delta G_{(\times.01)}\uparrow$ $E_{warp(\times 10^{-4})}\downarrow$ ProPainter (2023) 4.0 7.5 3.0 3.9 FGT (2022) 4.0 7.1 2.8 4.0 CN inpaint (2023a) 3.0 8.0 2.3 4.1 TokenFlow(2023) 2.0 6.9 1.4 5.7 RAVE (2023) 2.0 7.4 1.5 6.8 T2V-Zero (2023) 0.3 9.0 0.3 8.1 V-LASIK (ours) 4.0 7.6 3.0 3.6 V-LASIK masked (ours) 4.0 7.6 3.0 4.2
We compare the results of different video editing and inpainting methods to the results of our model in two versions — with and without masks. As mentioned in Sec. 4 both options exist for our model and there is a tradeoff between them. The masked version better preserves the original video colors, while the non-masked version better removes the glasses and thus is more temporally consistent. These differences are small, hence as shown in Tab. 1, they only slightly affect the quantitative results.
Fidelity: to test the fidelity of our results, we measure the average difference between the number of pixels with glasses in the original videos vs. the edited ones. To find the pixels that contain glasses in an image, we use a face parser (Zheng et al., 2022) that detects glasses, and apply it to each video frame. Then, we calculate the average difference between the number of pixels with glasses in the original frame and the edited one, normalized by the total number of pixels per frame, and report it as $\Delta G$ in Tab. 1. As the $\Delta G$ scores in Tab. 1 show, our method removes the glasses better than all other video editing methods, and is on-par with the inpainting methods ProPainter (Zhou et al., 2023) and FGT (Zhang et al., 2022). However, as illustrated in Figs. 5 and 6 and by the results of our user study in Fig. 7, although they remove the glasses from most videos, as our method does, they often generate unrealistic results with artifacts around the eyes, which our model does not generate.
Identity preservation: to test identity preservation, we use an ID score ($ID$ in Tab. 1), which is defined by the average cosine similarity between the face embeddings of the video frames, generated by the face recognition model Arcface (Deng et al., 2019). We note that neither of these metrics is complete on its own, as an unchanged video would get a perfect $ID$ score, and a random video without glasses would get a very high $\Delta G$ score. For example, as Text2Video-Zero (Khachatryan et al., 2023) does not remove the glasses from most videos, and does not change the videos by much, it achieves a very high $ID$ score. For this reason, we follow prior work (Kara et al., 2023; cong2023flatten) and also look at $ID\cdot\Delta G$, which quantifies the trade-off between removing the glasses from the video, and remaining faithful to the identity of the person in the original video. When looking at $ID\cdot\Delta G$, our model achieves the best results, together with ProPainter (Zhou et al., 2023). However, as mentioned, the results of our user-study in Fig. 7 show that our model outperforms ProPainter and all other methods both in terms of glasses-removal, and in terms of realism and identity preservation.
Temporal consistency: to evaluate temporal consistency, we follow previous work (Geyer et al., 2023; Zhou et al., 2023; Lei et al., 2020; Lai et al., 2018) and calculate the warp error ($E_{warp}$ in Tab. 1). We use RAFT (Teed and Deng, 2020) to calculate optical flow between each consecutive pair of frames, warp the former towards the latter, and calculate a masked MSE loss (masking occlusions) between the warped frame and the second frame. As shown in Tab. 1, the non-masked version of our model produces the most temporally consistent results, compared to all other methods.
5.4. Generalization
In this work, we focus on glasses as a case study. We chose glasses as they are a particularly complex task, as explained in Sec. 1. However, our method can be generalized to other types of local video editing.
5.4.1. Stickers
[…]
Figure 8. Stickers results. These examples show the generalization ability of our method. We apply it to removing facial stickers from videos, and show our model successfully removes different stickers from different locations over the face.
To show our method works for a different use-case, we generate a synthetic dataset of videos by applying stickers from the Stickers dataset (Nguyen et al., 2021) on faces from the CelebV-Text (Yu et al., 2023b) dataset, imitating real facial stickers, tattoos or synthetic features added by social-media apps, and show that our method is able to remove them. For this task, we detect the face in each frame, and estimate its 3D shape using a trained neural network that predicts the coefficients of the Basel 3D face model (Gerig et al., 2018). Then, we project and render the sticker texture on the original frame. For each generated clip, we sample one sticker and place it at a random position in the $uv$ coordinate space, which is mapped to the surface of the face. We generate 1274 clips this way, each 40-frames long.
After the dataset is created, we train the same image-to-image model over the stickers dataset and run our video editing pipeline using this model over the test videos. Some results are presented in Fig. 8. As shown, our model is able to seamlessly remove the stickers from the videos.
6. Limitations
Input Output Input Output No MM MM Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9. Limitations. Left: strong reflections. Middle: Dark sunglasses. In these cases, as the eyes are not exposed, the model cannot guess the right eye color. Moreover, our data was not clean and contained sunglasses videos which were treated the same way as eye glasses in the data generation process. Hence, the model generates dark areas instead of eyes for dark sunglasses videos. Right: Eye blurriness that comes from the motion module.
[…]
Although our method removes glasses reflections in many cases, it struggles with strong reflections, as in the left example in Fig. 9, where the true eyes are never visible. Additionally, our method struggles with dark sunglasses, as in the middle example in Fig. 9, where identifying eye color is difficult. Moreover, as current glasses detectors do not differ between eye and sun glasses, our dataset was not clean and contained sunglasses videos as well. As we used the same kind of masks for all our data, if the input is a video with sunglasses, our model often generates dark areas where the eyes should be, as is shown the middle example in Fig. 9. This could be solved by cleaning the data and only training over eye-glasses videos. Finally, the motion module tends to smooth the frames to get a more temporally consistent result. As a result, our outputs also tend to be a bit blurry, as shown in the right example in Fig. 9. As discussed in Sec. 4, by using only some of the motion layers, we reduce the blurriness to a minimum.
7. Societal impact
While the goal of this work is to allow editing of owned or licensed videos only, we acknowledge the fact it may be misused to altering videos without consent, and contribute to the spread of misinformation. We condemn such usage, and we are actively working on systems that detect synthetic and edited media.
8. Discussion
We explored the potential of local video editing through learning from imperfect synthetic data without paired data. Our results surpass existing methods, consistently and realistically removing glasses from videos while preserving the individual’s identity. We focus on the challenging case of removing glasses from videos, however we show our method can also be applied to other local video editing tasks such as removing stickers from faces. We hypothesize it would work similarly for any other local attribute, and we encourage future work to pursue this direction.
9. Acknowledgements
This project was funded in part by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme, grant agreement No. 802774 (iEXTRACT), and by ISF grant number 1337/22.